OpenStudy (anonymous):

Solve the systems using substitution. 3x + 2y = -4 x + 2y = -8

OpenStudy (anonymous):

x + 2y = -8 x = -8 - 2y 3( -8 - 2 y) + 2y = -4 -24 - 6y +2y = -4 -24 - 4y = -4 -4y = 20 y = -5 x = 2

OpenStudy (anonymous):

equation 1 : 3x + 2y = -4 equation 2 : x + 2y = -8 -> x = -2y -8 Let's plugin -2y-8 to first equation's x; 3(-2y-8) + 2y = -4 -6y - 24 + 2y = -4 -4y = 24-4 y= 20/ -4 y= -5 Plugin -5 to second equation's y; x = -2y -8 x=-2(-5)-8 x=10-8 x=2 Hope it helps.